OpenStudy (anonymous):

The vertices of a triangle are A(16, 0), B(9, 2), and C(0, 0). Which side is also an altitude with a slope of -2/7?

OpenStudy (anonymous):

ok , so u have 3 coordinate points, use the formula:\[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] to get the slope for each line. And then compare ur answer!

OpenStudy (anonymous):

so you mean get the slope of A&B then B&C and C&A then compare?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

line 1 = AB line 2 = BC line 3 = CA

OpenStudy (anonymous):

A&B = -2/7 B&C = 9/2 C&A = 0/-16

OpenStudy (anonymous):

so its A&B

OpenStudy (anonymous):

tht is right!